# Many-body non-Hermitian Skin Effect for Multipoles

### v1

@@ Abstract
Figure 3: Dynamical probes of the late-time steady-state after evolution with a charge-conserving (inset: dipole-conserving) non-Hermitian Hamiltonian with non-reciprocal hopping imbalance $g$. (a) The half-chain von Neumann entanglement entropy exhibits area-law scaling. (b) The inverse participation ratio converges with system size, indicating Fock space localization. Both quantities are averaged over random initial product states.
Dynamical signatures.—Having defined and examined the skin effect in systems that conserve multipole moments, we now focus on its implications for dynamics. Dynamical probes are especially important because they can be implemented in experimental platforms to characterize the skin effect. Concretely, we will study the state of a system at time $t$ undergoing dynamics modeled by a non-Hermitian Hamiltonian $H$:
$|\psi(t)\rangle=\frac{e^{-itH}|\psi(0)\rangle}{\|e^{-itH}|\psi(0)\rangle\|}.$ (9)
We have already discussed the dynamics of charge: the $m$-pole skin effect causes charged particles to flow and generate an $(m+1)$-pole moment. Now, we turn to the evolution of quantum correlations, focusing on the entanglement scaling. Entanglement is quantified by the von Neumann entropy $S(t)=-\text{Tr}[\rho_{L/2}\log{\rho_{L/2}}]$, where $\rho_{L/2}$ is the bipartite reduced density matrix of the evolved state $|\psi(t)\rangle$ for a spatial cut the divides the system in half. Generic closed quantum systems, such as $H_{\text{dip}}(0,0)$, thermalize under their own time evolution. It follows that generic states evolve to have volume-law entanglement, $S\sim{O}(L)$. This behavior should be distinguished from area-law entanglement, $S\sim O(1)$, which is typically found only for ground states.
In addition to entanglement, we also investigate the spread of $|\psi(t)\rangle$ in Fock space through the inverse participation ratio, $IPR_{2}(t)=\sum_{\overline{n}}|\langle\overline{n}|\psi(t)\rangle|^{4}$, where $\{|\overline{n}\rangle\}$ is the many-body computational basis. The evolved state spans the entire Hilbert space if $IPR_{2}\sim 2^{-D_{2}L}$ with fractal dimension $D_{2}=1$, and is localized if $D_{2}=0$. Let us consider a generic $m$-pole conserving Hamiltonian having asymmetric hopping such that a similarity transformation $R^{-1}H^{m}(g)R=H^{m}(0)$ exists, with $R=e^{-\beta(g)P^{(m+1)}}=e^{-\beta(g)\sum_{j}j^{m+1}\rho_{j}}$. Starting from a random product state in the computational basis, |$\ket{\psi(0)}=\prod_{\{s_{j}\}}c_{s_{j}}^{\dagger}\ket{0}$, we can use the transformation to write
$IPR_{2}(t)=\frac{\sum_{\underline{n}}e^{-4\beta\sum_{j}j^{m+1}n_{j}}|\langle\underline{n}|e^{-itH^{(m)}(0)}|\psi(0)\rangle|^{4}}{(\sum_{\underline{n}}e^{-2\beta\sum_{j}j^{m+1}n_{j}}|\langle\underline{n}|e^{-itH^{(m)}(0)}|\psi(0)\rangle|^{2})^{2}}.$ (10)
@@
$|{IPR_{2}(t\rightarrow\infty)}|\approx\prod_{j=1}^{L}\frac{(1+e^{-4\beta j^{m+1}})}{(1+e^{-2\beta j^{m+1}})^{2}}.$ (11)
@@
This expression is bounded by |$\log{IPR_{2}}|\leq\sum_{j=1}^{L}\left|\log{\left(1-\frac{2e^{-2\beta j^{m+1}}}{(1+e^{-2\beta j^{m+1}})^{2}}\right)}\right|$ ) |, which converges as $L\rightarrow\infty$, implying localization in Fock space ($D_{2}\rightarrow 0$). This should be viewed as a generalization of the edge localization that results from the single-particle skin effect. Although fermion repulsion prevents many-body states from localizing at the physical boundary, the $m$-pole skin effect extremizes the $(m+1)$-pole moment of any initial state, driving time evolution to a highly restricted region of the Fock space. Finally, localization in Fock space implies that the entanglement entropy of the steady-state has area-law scaling [71].
@@
We test our conjecture by performing exact numerical simulations of charge- and dipole-conserving time evolution. Fig. 3 (a-b) shows $S(t\rightarrow\infty)$ and $IPR_{2}(t\rightarrow\infty)$ for the $U(1)$ charge skin effect using $H_{\text{HN}}(g)$ in Eq. 1 (for various system sizes as a function of the non-reciprocity $g$). In the insets of Fig. 3 (a-b), we consider the case of the dipole skin effect with $H_{4}(g)$ in Eq. 5. In agreement with our analytical analysis, we find that the final states exhibit area-law scaling for the entanglement entropy and are localized in Fock space.
@@
Conclusions— Nonreciprocal hopping leads to the localization of charge at one boundary of a 1D chain. On the other hand, $m$-pole conservation imposes strong constraints on charge transport. The competition between these two effects gives rise to a new class of non-Hermitian skin effects in intrinsically interacting systems. In this work, we studied these generalized skin effects using lattice models, field theory arguments, and numerics. Our central result, that eigenstates exhibiting the $m$-pole skin effect have an $(m+1)$-pole moment, naturally implies a definition of the ordinary skin effect beyond the non-interacting regime.
@@
Furthermore, this definition provides an experimentally accessible signature of the $m$-pole skin effect, which may then be observed in ultra-cold atom platforms. For example, non-reciprocal charge hopping can be engineered by coupling an optical lattice to a collective one-body loss term induced by dissipation [72]. Placing this lattice in a strongly tilted external field then effectively imposes dipole-conservation [73, 74, 75], and the ensuing dynamics is governed by the dipole skin effect. Looking beyond local densities of conserved charges, we have also provided a truly quantum characterization of the skin effect, showing that it results in area-law scaling of the entanglement entropy in long-time steady-states.
@@
We have seen an intimate connection between the symmetries of a system and its non-Hermitian dynamics. To this end, it would be interesting to consider models with other types of symmetries, like higher-form or subsystem symmetries, as well as patterns of symmetry-breaking. It may also be possible to extend the $m$-pole skin effect to higher dimensions and connect it to many-body topological invariants, both fruitful directions for future research.
@@
G. D. T. acknowledges the support from the EPiQS Program of the Gordon and Betty Moore Foundation. T.L.H. and J. G. acknowledge support from the U.S. Office of Naval Research (ONR) Multidisciplinary University Research Initiative (MURI) under Grant No. N00014-20-1-2325 on Robust Photonic Materials with Higher-Order Topological Protection.

### v2

@@ Abstract
Figure 3: Dynamical probes of the late-time steady-state after evolution with charge-conserving (inset: dipole-conserving) non-Hermitian Hamiltonian with non-reciprocal hopping imbalance $g$, specifically $H_{\text{HN}}(g)$ ($H_{\text{dip}}(g,3g/2)$). (a) The half-chain von Neumann entanglement entropy exhibits area-law scaling. (b) The inverse participation ratio converges with system size, indicating Fock-space localization. Both quantities are averaged over random initial product states.
Dynamical signatures.—Having defined and examined the skin effect in systems that conserve multipole moments, we now focus on its implications for dynamics. Dynamical probes are especially important because they can be implemented in experimental platforms to characterize the skin effect. Concretely, we study the state of a system at time $t$ undergoing dynamics modeled by a non-Hermitian Hamiltonian $H$:
$\begin{gathered}|\psi(t)\rangle=\frac{e^{-itH}|\psi(0)\rangle}{\|e^{-itH}|\psi(0)\rangle\|}.\end{gathered}$ (9)
We have already discussed the dynamics of charge: the $m$-pole skin effect causes charged particles to flow and generate an $(m+1)$-pole moment. Now, we turn to the evolution of quantum correlations, focusing on the spread of entanglement. Entanglement is quantified by the von Neumann entropy $S(t)=-\text{Tr}[\rho_{L/2}\log{\rho_{L/2}}]$, where $\rho_{L/2}$ is the reduced density matrix of the evolved state $|\psi(t)\rangle$ for a spatial cut that divides the system in half. Generic closed quantum systems, such as $H_{\text{dip}}(0,0)$, thermalize under their own time evolution. It follows that generic states evolve to have volume-law entanglement, $S\sim{O}(L)$. This behavior should be distinguished from area-law entanglement, $S\sim O(1)$, which is typically found only for ground states.
Besides entanglement, we also investigate the spread of $|\psi(t)\rangle$ in Fock space through the inverse participation ratio, $IPR_{2}(t)=\sum_{\overline{n}}|\langle\overline{n}|\psi(t)\rangle|^{4}$, where $\{|\overline{n}\rangle\}$ is the many-body computational basis. The evolved state spans the entire Hilbert space if $IPR_{2}\sim 2^{-D_{2}L}$ with fractal dimension $D_{2}=1$, and is localized if $D_{2}=0$. Let us consider a generic $m$-pole conserving Hamiltonian with asymmetric hopping such that a similarity transformation $R^{-1}H^{m}(g)R=H^{m}(0)$ exists, with $R=e^{-\beta(g)P^{(m+1)}}=e^{-\beta(g)\sum_{j}j^{m+1}\rho_{j}}$. Starting from a random product state in the computational basis, |$\ket{\psi(0)}=\prod_{\{s_{j}\}}c_{s_{j}}^{\dagger}\ket{0}$, we can use the transformation to write
$\begin{gathered}IPR_{2}(t)=\frac{\sum_{\underline{n}}e^{-4\beta\sum_{j}j^{m+1}n_{j}}|\langle\underline{n}|e^{-itH^{(m)}(0)}|\psi(0)\rangle|^{4}}{(\sum_{\underline{n}}e^{-2\beta\sum_{j}j^{m+1}n_{j}}|\langle\underline{n}|e^{-itH^{(m)}(0)}|\psi(0)\rangle|^{2})^{2}}.\end{gathered}$ (10)
@@
$\begin{gathered}|{IPR_{2}(t\rightarrow\infty)}|\approx\prod_{j=1}^{L}\frac{(1+e^{-4\beta j^{m+1}})}{(1+e^{-2\beta j^{m+1}})^{2}}.\end{gathered}$ (11)
@@
This expression is bounded by |$\log{IPR_{2}}|\leq\sum_{j=1}^{L}\left|\log{\left(1-\frac{2e^{-2\beta j^{m+1}}}{(1+e^{-2\beta j^{m+1}})^{2}}\right)}\right|$ ) |, which converges as $L\rightarrow\infty$, implying localization in Fock space ($D_{2}\rightarrow 0$). This should be viewed as a generalization of the edge localization that results from the single-particle skin effect  [^10] The strength of the localization can also be seen using the similarity transformation, which breaks up the evolution into a unitary piece, which generically yields a completely delocalized state, and the similarity map, $e^{-\beta(g)P^{(m+1)}}$, which then localizes this state in the corner of the Fock space with maximal $(m+1)$-pole moment. Consequently, $\beta(g)\sim g$ can be interpreted as a localization length. . Although fermion repulsion prevents many-body states from localizing at the physical boundary, the $m$-pole skin effect extremizes the $(m+1)$-pole moment of any initial state, driving time evolution to a highly restricted region of the Fock space. Finally, Fock-space localization implies that the entanglement entropy of the steady-state has area-law scaling De Tomasi and Khaymovich (2020).
@@
We test our conjecture by performing exact numerical simulations of charge- and dipole-conserving time evolution using free fermion techniques for the non-interacting case and Chebyshev polynomial methods for the interacting case Weiße et al. (2006); Bera et al. (2017). Fig. 3(a-b) shows $S(t\rightarrow\infty)$ and $IPR_{2}(t\rightarrow\infty)$ for the $\text{U}(1)$ charge skin effect using $H_{\text{HN}}(g)$ in Eq. (1) (as a function of non-reciprocity $g$ and for various system sizes). In the insets of Fig. 3(a-b), we consider the dipole skin effect with $H_{4}(g)$ in Eq. (5). In agreement with our analysis, we find that the final states exhibit area-law scaling for the entanglement entropy and are localized in Fock space.
@@
Conclusions— Non-reciprocal hopping leads to the localization of charge at one boundary of a 1D system. On the other hand, $m$-pole conservation imposes strong constraints on charge transport. The competition between these two effects gives rise to a new class of non-Hermitian skin effects in intrinsically interacting systems. In this work, we studied these generalized skin effects using lattice models, field theory arguments, and numerics. Our central result, that eigenstates exhibiting the $m$-pole skin effect have an $(m+1)$-pole moment, naturally implies a definition of the ordinary skin effect beyond the non-interacting regime.
@@
Furthermore, this definition provides an experimentally accessible signature of the $m$-pole skin effect, which may then be observed in ultra-cold atom platforms. For example, non-reciprocal charge hopping can be engineered by coupling an optical lattice to external dissipation Gong et al. (2018). Placing this lattice in a strongly tilted external field then effectively imposes dipole-conservation Guardado-Sanchez et al. (2020); Scherg et al. (2021); Kohlert et al. (2023), and the ensuing dynamics is governed by the dipole skin effect. Looking beyond local densities of conserved charges, we have also provided a truly quantum characterization of the skin effect, showing that it results in area-law entanglement entropy in steady-states.
@@
We have seen an intimate connection between the symmetries of a system and its non-Hermitian dynamics. To this end, it would be interesting to consider models with other types of symmetries, like spinful models with time-reversal symmetry, or higher-form and subsystem symmetries, as well as patterns of symmetry-breaking. It may also be possible to extend the $m$-pole skin effect to higher dimensions and connect it to many-body topological invariants, both fruitful directions for future research.
Acknowledgements.
@@
G.D.T. acknowledges the support from the EPiQS Program of the Gordon and Betty Moore Foundation. T.L.H. and J.G. acknowledge support from the U.S. Office of Naval Research (ONR) Multidisciplinary University Research Initiative (MURI) under Grant No. N00014-20-1-2325 on Robust Photonic Materials with Higher-Order Topological Protection. G.D.T. acknowledges the hospitality of MPIPKS Dresden where part of the work was done.
Supplementary Material
SM 1: Dipole skin effect in the absence of a similarity transformation to a Hermitian system
In the main text, much of our investigation of the dipole skin effect (and its higher multipole relatives) has been limited to non-Hermitian systems with a real spectrum in open boundary conditions (OBCs). In these cases, there exists a similarity transformation that maps the non-Hermitian Hamiltonian to a Hermitian one, greatly facilitating the analysis. Nevertheless, we heuristically argued using a response field theory that the dipole skin effect is also present in more general non-Hermitian Hamiltonians with a complex spectrum. Here, we provide some further numerical evidence for this claim.
Figure S1: Energies in OBCs ($L=12$) for $H_{\text{dip}}(g_{1},g_{2})$ for cases where a similarity transformation exists (red dots, $g_{1}=2/3,g_{2}=1$) and does not exist (blue dots, $g_{1}=2,g_{2}=1$). The former are always real, while the latter are complex.
We focus on the dipole-conserving non-reciprocal Hamiltonian that we introduced in Eq. (7) of the main text, which we reproduce here for convenience:
$H_{\text{dip}}(g_{1},g_{2})$ $=-\sum_{j}\left(h^{(2)}_{j}(g_{1})+t\,h^{(3)}_{j}(g_{2})\right),$ (S1)
[…]
We henceforth fix $t=1$, and work in the symmetry sector with zero dipole moment and magnetization. For $g_{1}=2g_{2}/3$, this Hamiltonian has a real energies in open boundary conditions. Consequently, the similarity transformation performed by $R=e^{-gP^{(2)}/4}$, where $P^{(2)}$ is the total quadrupole moment, maps this Hamiltonian to a Hermitian model.
For $g_{1}=2g_{2}$, on the other hand, the OBC spectrum is non longer real. Nevertheless, the key features of the dipole skin effect are the same as the model explored in the main text. In Fig. S1, we show the spectrum of this system for an open chain of length $L=12$. In Fig. S2, we plot the local charge distribution for 40 mid-spectrum OBC eigenstates of this Hamiltonian for $L=20$, all of which have an extensive quadrupole moment. Lastly, the charge distribution after time evolving a random initial state with $H_{\text{dip}}(2g,g)$ is shown in Fig. S3. As expected, the steady-state displays a symmetrical pattern of charge that reflects a maximal quadrupole moment.
Figure S2: Local charge distributions, $\rho_{j}$, for 40 mid-spectrum eigenstates of $H_{\text{dip}}(g_{1},g_{2})$ with $g_{1}=2,g_{2}=1$, for an open chain of length $L=20$. All eigenstates exhibit a charge distribution characteristic of a near-maximal quadrupole moment.
Figure S3: Time evolution of a random state (averaged over 20 trials) using $H_{\text{dip}}(g_{1},g_{2})$ with $g_{1}=2,g_{2}=1$. The steady state exhibits a quadrupole moment of the charge.
SM 2: Dipole skin effect in the presence of dipole-symmetry breaking terms
Starting from a dipole-conserving Hamiltonian such as Eq. (S1) and then adding charge hopping terms, as in the Hatano-Nelson model, explicitly breaks dipole conservation. As a result, it does not make sense to project the Hamiltonian into a sector of fixed dipole moment and look at the distribution of eigenstates. While one can still project into a sector of fixed charge (as it remains conserved), the resulting eigenstates will have very different charge distributions depending on their different dipole moments, obscuring the skin effect.
However, the dipole skin effect can still be seen in dynamics. In particular, we take a Hamiltonian with both dipole-hopping and charge-hopping terms:
$H=H_{\text{dip}}(g_{1},g_{2})-\delta\sum_{j}(e^{g}c^{\dagger}_{j}c_{j+1}+e^{-g}c^{\dagger}_{j+1}c_{j}),$ (S2)
where the second term is simply the Hatano-Nelson model, and $\delta$ controls the strength of dipole symmetry-breaking. When $\delta$ is smaller than the other scales in the system, the dipole skin effect continues to approximately hold, and a random initial state evolves under Eq. (S2) to a state with a clear quadrupole moment. As $\delta$ becomes large and the charge-hopping dominates, the steady state instead exhibits a dipole moment, a signature of the charge skin effect dominating over the dipole skin effect.
Figure S4: Time evolution of a random state using a Hamiltonian with both charge and dipole hopping terms (Eq. (S2) with $g_{1}=g_{2}=1$ and $g=1.5$), where $\delta$ controls the relative strength of charge-hopping over dipole hopping. (a) With small enough breaking of dipole symmetry ($\delta=0.1$), the dipole skin effect approximately holds, (b) large dipole symmetry-breaking ($\delta=10$) yields the charge skin effect.
These two scenarios are shown in Fig. S4.